OpenStudy (anonymous):

Write a polynomial function f of least degree that has rational coefficients, a leading coefficient of 1, and the given zeros.

OpenStudy (cwrw238):

where are the given zeros?

OpenStudy (anonymous):

yes -2,-1,1

OpenStudy (cwrw238):

f(x) = (x + 2)(x + 1)(x - 1) expand this to get the polynomial

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

how do you expand it ?

OpenStudy (cwrw238):

(x + 1)(x - 1) = x(x - 1) + 1(x - 1) = x^2 - x + x - 1 = x^2 - 1 now work to (x + 2)(x^2 - 1) in the same way

OpenStudy (anonymous):

thank you so much

OpenStudy (cwrw238):

yw